# Dynamical Dark Energy from Lattice Quantum Gravity

### v2

@@ III.4 Determining Newton’s constant and setting the lattice spacing
The results for $G/a_{\rm lat}^{2}$ in the infinite-volume limit, denoted by $G_{a}$ in the fit function Eq. (39), are given in Tab. 6 for all four lattice spacings. Taking $\kappa_{2}=3.0$ as our fiducial lattice spacing, the values of $a_{\rm rel}$ can be obtained from these results; these are also shown in Tab. 6. We see that as $\kappa_{2}$ increases, $a_{\rm rel}$ decreases, so that larger $\kappa_{2}$ corresponds to finer lattice spacing, as expected. Knowing the dimensionless ratio $G/a_{\rm lat}^{2}$ allows us to convert $a_{\rm lat}$ into units of the Planck length, where $G/a_{\rm lat}^{2}>1$ would indicate that $a_{\rm lat}$ is sub-Planckian. Thus, we see that our finest lattice spacing measured in $a_{\rm lat}$ units is very nearly the Planck length, though there appears to be no limitation other than practical for pushing it smaller. Using our $a_{\rm lat}/\ell$ values from Sec. III.3 to express our lattice spacings in dual lattice units, and again taking $\kappa_{2}=3.0$ as the fiducial spacing so that $\ell_{\rm rel}=1$ there, we find the values of $\ell_{\rm rel}$ given in Tab. 6. Note that our results for $G/a_{\rm lat}^{2}$, combined with our results for $a_{\rm lat}/\ell$, imply that $\ell$ is significantly smaller than the Planck length. The dual length $\ell$ is around 23 times smaller than the Planck length on our finest lattice spacing. The relative dual length $\ell_{\rm rel}$ also decreases as $\kappa_{2}$ is taken larger, further confirming our picture that a continuum limit can be taken by following the first order line in the phase diagram of Figure 1 out to large $\kappa_{2}$ values.
@@ IV.1 Theory and motivation
Our lattice geometries are well-described by the semi-classical Euclidean de Sitter solution, which is solid evidence that EDT could define a non-perturbative formulation of quantum gravity. Interestingly, a more careful study shows that there are corrections to this picture. As pointed out in Section III.3.2, we find that the cosmological constant is not strictly constant as a function of Euclidean time, particularly on our smaller volume ensembles. A deviation from the expected classical scaling of the geometries with volume could provide a window to quantum gravitational effects, assuming that they survive a continuum limit. Although such deviations might be treated as finite-size effects to be extrapolated away in a flat space lattice QCD calculation, in quantum gravity the dynamics of the universe at small sizes is an interesting question in its own right. Figure 12 shows the cosmological constant inferred from Eq. 36 as a function of Euclidean time $\tau$ on one of our smaller physical-volume ensembles. Eq. 36 requires as input the infinite-volume limit of $a_{\rm lat}/\ell$, which is needed to set $\ell_{V}/\ell$. Although all of our ensembles are consistent with the classical de Sitter solution, at least within a chosen fit window, this implies a particular finite-volume value of $a_{\rm lat}/\ell$. Because the lattice volumes at which we are simulating are rather small in Planck units, and the classical de Sitter solution that is used to fix $a_{\rm lat}/\ell$ only applies at large volumes, we must extrapolate our values of $a_{\rm lat}/\ell$ to the infinite-volume limit before we can use them as conversion factors. When we use the appropriate infinite-volume extrapolated values of $a_{\rm lat}/\ell$ in Eq.  36, the behavior of $\Lambda$ versus $\tau$ on a finite volume shows significant deviations from the pure de Sitter solution, where a constant $\Lambda$ is expected. We find that these results are broadly consistent across lattice spacings and that the Euclidean time dependence of the effective cosmological constant does not appear to be a lattice artifact. In the remainder of this work we show that these deviations can be described by nontrivial vacuum dynamics and that they pass a number of important consistency checks in this interpretation.
@@
for some constant $b$. We find Eq. 40 is a good description of the lattice data if we identify the scale $\mu$ with the Hubble rate $H=\dot{A}/A$. The absolute value in Eq. (40) reflects the fact that the renormalization scale should be identified with the magnitude of the Hubble rate, and that the $\mu^{2}$ term does not pick up a minus sign due to the Euclidean continuation of the time derivative appearing in $H$. Note that in our simulations $H$ is not a constant as a function of Euclidean time because the spatial curvature term is not negligible, as it would be for pure de Sitter space in Lorentzian signature at large times. Figure 13 (bottom-right panel) shows the same data for $\Lambda(\tau)$ as in Fig. 12, now plotted as a function of $H^{2}$, which is reconstructed from the first derivative of the scale factor $A(\tau)$. This set of data is well-described by a fit linear in $H^{2}$, hinting that our results could be consistent with a picture in which the vacuum energy runs according to Eq. 40, with a renormalization scale associated with the Hubble rate. A more comprehensive study incorporating multiple lattice spacings and volumes shows that our results are well-described by this picture, allowing a first principles determination of the parameters $\nu$ and $b$.
Figure 12: $\Lambda$ versus $\tau$ at our finest lattice spacing, with $\kappa_{2}=3.8$ and $N_{4}=16000$.
@@ IV.2 Determining model parameters from the lattice
The calculation of the parameters of the running vacuum model $\nu$ and $\tilde{\nu}$ in Eq. (42) using our lattice results proceeds as follows. The scale factor measured in simplex units is obtained directly from the lattice scale factor from
$a(\tau)=\frac{\ell_{V}}{\ell}A(\tau),$ (56)
where we use the notation $A(\tau)\equiv(N_{4}^{\rm shell}(\tau))^{\frac{1}{3}}$ introduced in Sec. III.3.2 to emphasize the dependence on the lattice spacing conversion factors. The factor $\ell_{V}/\ell$ is obtained from the results for $a_{\rm lat}/\ell$ from our combined analysis in Section III using
The parameters of the dark energy model can then be constructed from the scale factor $A(\tau)$ and its derivatives with respect to $\tau$, including the Hubble rate $H=\dot{A}/A$, and the quantities $S=\ddot{A}/A$ and $T=\dddot{A}/\dot{A}$, where all three of these are measured in dual lattice units.
@@
and the derivatives are all taken under a jackknife to correctly propagate the errors through to the next stage of the analysis.
@@
$3\left(-\frac{\dot{A}^{2}}{A^{2}}+\frac{\ell^{2}}{\ell_{V}^{2}}\frac{1}{A^{2}}\right)=\Lambda(H)=I_{H^{2}}+3\nu^{\prime}H^{2}.$ (58)
The sign difference between the two terms on the left side of Eq. (58) leads to a substantial cancellation between these terms. In order to construct this combination with sufficient precision to extract $\nu^{\prime}$, we require a precise determination of $\ell_{V}/\ell$ (which itself depends on having a precise $a_{\rm lat}/\ell$). This partial cancellation sets the precision level needed for the $a_{\rm lat}/\ell$ determination of Sect. III, which must be at the percent level in order to resolve a $\nu^{\prime}$ statistically and systematically different from zero. The value of $\nu^{\prime}$ is obtained from the slope of a linear fit to the left side of Eq. (58) as a function of $H^{2}$, and $I_{H^{2}}\equiv\Lambda_{0}+3\tilde{\nu}S$ is obtained from the intercept. As we show in the following subsections, $\tilde{\nu}$ is consistent with zero. Taking $\tilde{\nu}=0$, we have $I_{H^{2}}=\Lambda_{0}$, and this becomes the input to our Method 2 determination of $a_{\rm lat}/\ell$, as detailed in Sect. III.3.2.
@@ IV.3.1 Fits to $\Lambda(H)$ versus $H^{2}$
As discussed in the previous subsection, the value of $\nu^{\prime}$ is obtained by using the lattice shelling function to construct the left side of Eq. (58). This leads to an effective cosmological constant $\Lambda(H)$, which must be linear in $H^{2}$ if it is to be consistent with Eq. (42). Fig. 13 shows the effective $\Lambda$ plotted versus $H^{2}$ for representative ensembles across four lattice spacings, where we find solid evidence that $\Lambda(H)$ is linear in $H^{2}$. The linear fit region that we chose is marked with black points in Fig. 13.
@@
Figure 13: $\Lambda$ versus $H^{2}$ plotted for four lattice spacings, along with a linear fit to the form Eq. (60). The points included in the fits are highlighted in black. Top left panel: $N_{4}=6$k, $\kappa_{2}=2.45$. Top right panel: $N_{4}=16$k, $\kappa_{2}=3.0$. Bottom left panel: $N_{4}=16$k, $\kappa_{2}=3.4$. Bottom right panel: $N_{4}=16$k, $\kappa_{2}=3.8$.
@@ IV.3.2 Fits to $\nu^{\prime}$ as a function of four-volume
Our synthetic data sample is obtained from 10,000 random Gaussian draws for every ensemble, with a mean given by the central value of $\nu^{\prime}$ and the error given by the statistical and fit systematic error added in quadrature. We account for the error coming from $a_{\rm lat}/\ell$ to the arrays as follows: We assume that when $a_{\rm lat}/\ell$ for a given lattice spacing deviates from its central value by some percentage of one $\sigma_{a_{\rm lat}/\ell}$, all $\nu^{\prime}$ values at the same lattice spacing also shift by the same percentage of its $a_{\rm lat}/\ell$ systematic error, $\sigma_{\nu^{\prime},a_{\rm lat}/\ell}$. Therefore, for each lattice spacing, we generate an array of $10,000$ samples from a Gaussian random number generator, $A_{\rm ran}\thicksim\mathcal{N}(0,1)$. This allows us to construct the array $\left(\frac{a_{\rm lat}}{\ell}\right)_{\rm avg}+A_{\rm ran}\sigma_{\frac{a_{\rm lat}}{\ell}}$, which gives a distribution for $\frac{a_{\rm lat}}{\ell}$ with the correct central value and standard deviation at each lattice spacing. To obtain the final correlated array for $\nu^{\prime}$ we use the fact that for two independent normal distributions, $\mathcal{N}_{1}(\mu_{1},\sigma_{1})$, $\mathcal{N}_{2}(\mu_{2},\sigma_{2})$, their sum $\mathcal{N}_{1}+\mathcal{N}_{2}=\mathcal{N}(\mu_{1}+\mu_{2},\sqrt{\sigma_{1}^{2}+\sigma_{2}^{2}})$. We then get $\nu^{\prime}$ for each ensemble by subtracting $A_{\rm ran}\sigma_{\nu^{\prime},\frac{a_{\rm lat}}{\ell}}$ from the original $\nu^{\prime}$ sample thereby including the $a_{\rm lat}/\ell$ systematic error in quadrature with the statistical and fit systematic error in $\nu^{\prime}$, and maintaining the correlation with $a_{\rm lat}/\ell$. The minus sign is due to the fact that the larger $a_{\rm lat}/\ell$ is, the smaller $\nu^{\prime}$ is. Because the random array for $a_{\rm lat}/\ell$ is shared by all ensembles at the same lattice spacing, the $a_{\rm lat}/\ell$ error is fully correlated across different volumes at a single lattice spacing. The fully correlated $a_{\rm lat}/\ell$ error covariance matrix is added to the diagonal matrix containing the statistical and fitting errors of $\nu^{\prime}$ in quadrature. This gives the full covariance matrix for $\nu^{\prime}$, which is reused when performing the fits over a jackknife array of $10,000$ random samples.
@@ IV.5 Alternative calculation of $\nu^{\prime}$
Unfortunately, the third derivative $\dddot{A}$ is only poorly determined by the data, given the large statistical errors associated with numerical derivatives of high order. There is also expected to be significant cancellation between the $\tilde{\nu}$ term proportional to $S$ with that proportional to $T$, since $S$ and $T$ are equal up to corrections of ${\cal O}(H^{2})$, so that a meaningful constraint on $\tilde{\nu}$ is not possible with this approach.

### v3

@@ III.4 Determining Newton’s constant and setting the lattice spacing
The results for $G/a_{\rm lat}^{2}$ in the infinite-volume limit, denoted by $G_{a}$ in the fit function Eq. (42), are given in Tab. 6 for all four lattice spacings. Taking $\kappa_{2}=3.0$ as our fiducial lattice spacing, the values of $a_{\rm rel}$ can be obtained from these results; these are also shown in Tab. 6. We see that as $\kappa_{2}$ increases, $a_{\rm rel}$ decreases, so that larger $\kappa_{2}$ corresponds to finer lattice spacing, as expected. Knowing the dimensionless ratio $G/a_{\rm lat}^{2}$ allows us to convert $a_{\rm lat}$ into units of the Planck length, where $G/a_{\rm lat}^{2}>1$ would indicate that $a_{\rm lat}$ is sub-Planckian. Thus, we see that our finest lattice spacing measured in $a_{\rm lat}$ units is very nearly the Planck length, though there appears to be no limitation other than practical for pushing it smaller. Using our $a_{\rm lat}/\ell$ values from Sec. III.3 to express our lattice spacings in renormalized dual lattice units, and again taking $\kappa_{2}=3.0$ as the fiducial spacing so that $\ell_{\rm rel}=1$ there, we find the values of $\ell_{\rm rel}$ given in Tab. 6. Note that our results for $G/a_{\rm lat}^{2}$, combined with our results for $a_{\rm lat}/\ell$, imply that $\ell$ is significantly smaller than the Planck length. The length $\ell$ is around 23 times smaller than the Planck length on our finest lattice spacing. The relative length $\ell_{\rm rel}$ also decreases as $\kappa_{2}$ is taken larger, further confirming our picture that a continuum limit can be taken by following the first order line in the phase diagram of Figure 1 out to large $\kappa_{2}$ values.
@@ IV.1 Theory and motivation
Our lattice geometries are well-described by the semi-classical Euclidean de Sitter solution, which is solid evidence that EDT could define a non-perturbative formulation of quantum gravity. Interestingly, a more careful study shows that there are corrections to this picture. As pointed out in Section III.3.2, we find that the cosmological constant is not strictly constant as a function of Euclidean time, particularly on our smaller volume ensembles. A deviation from the expected classical scaling of the geometries with volume could provide a window to quantum gravitational effects, assuming that they survive a continuum limit. Although such deviations might be treated as finite-size effects to be extrapolated away in a flat space lattice QCD calculation, in quantum gravity the dynamics of the universe at small sizes is an interesting question in its own right. Figure 12 shows the cosmological constant inferred from Eq. 38 as a function of Euclidean time $\tau/\ell$ on one of our smaller physical-volume ensembles. Eq. 38 requires as input the infinite-volume limit of $a_{\rm lat}/\ell$, which is needed to set $\ell_{V}/\ell$. Although all of our ensembles are consistent with the classical de Sitter solution, at least within a chosen fit window, this implies a particular finite-volume value of $a_{\rm lat}/\ell$. Because the lattice volumes at which we are simulating are rather small in Planck units, and the classical de Sitter solution that is used to fix $a_{\rm lat}/\ell$ only applies at large volumes, we must extrapolate our values of $a_{\rm lat}/\ell$ to the infinite-volume limit before we can use them as conversion factors. When we use the appropriate infinite-volume extrapolated values of $a_{\rm lat}/\ell$ in Eq.  38, the behavior of $\Lambda$ versus $\tau$ on a finite volume shows significant deviations from the pure de Sitter solution, where a constant $\Lambda$ is expected. We find that these results are broadly consistent across lattice spacings and that the Euclidean time dependence of the effective cosmological constant does not appear to be a lattice artifact. In the remainder of this work we show that these deviations can be described by nontrivial vacuum dynamics and that they pass a number of important consistency checks in this interpretation.
@@
for some constant $b$. We find Eq. 43 is a good description of the lattice data if we identify the scale $\mu$ with the Hubble rate $H=\dot{A}/A$, or at the largest scales probed on a given ensemble, by the four-volume. At these largest scales the presence of the classical maximum of the shelling function ensures that $H=0$ at that point. Even so, we find that $\nu$ does not go to zero there, as Eq. (44) would imply if the scale $\mu$ were to be associated with the Hubble rate $H$ and set to zero. In this case the (inverse) four volume serves as an alternative cut-off in Eq. (44). The absolute value in Eq. (43) reflects the fact that the renormalization scale should be identified with the magnitude of the Hubble rate, and that the $\mu^{2}$ term does not pick up a minus sign due to the Euclidean continuation of the time derivative appearing in $H$. Note that in our simulations $H$ is not a constant as a function of Euclidean time because the spatial curvature term is not negligible, as it would be for pure de Sitter space in Lorentzian signature at large times. Figure 13 (bottom-right panel) shows the same data for $\Lambda(\tau)$ as in Fig. 12, now plotted as a function of $H^{2}$, which is reconstructed from the first derivative of the scale factor $A(\tau)$. This set of data is well-described by a fit linear in $H^{2}$, hinting that our results could be consistent with a picture in which the vacuum energy runs according to Eq. 43, with a renormalization scale associated with the Hubble rate. A more comprehensive study incorporating multiple lattice spacings and volumes shows that our results are well-described by this picture, allowing a first principles determination of the parameters $\nu$ and $b$.
Figure 12: $\Lambda$ versus $\tau/\ell$ at our finest lattice spacing, with $\kappa_{2}=3.8$, $\beta=0.92$ and $N_{4}=16000$. The points in black are the same ones used in the fit appearing in the lower right frame of Fig. 13. The inflection point discussed in III.3.1 and the peak of the shelling function are located at $\tau/\ell=43$ and $\tau/\ell=71$, respectively. These points are highlighted in red.
@@ IV.2 Determining model parameters from the lattice
The calculation of the parameters of the running vacuum model $\nu$ and $\tilde{\nu}$ in Eq. (45) using our lattice results proceeds as follows. The scale factor measured in $\ell$ units is obtained directly from the lattice scale factor from
$\frac{a(\tau)}{\ell}=\frac{\ell_{V}}{\ell}\frac{A(\tau)}{\ell_{V}},$ (59)
where we use the notation $A(\tau)/\ell_{V}\equiv(N_{4}^{\rm shell}(\tau))^{\frac{1}{3}}$ introduced in Sec. III.3.2 to emphasize the dependence on the lattice spacing conversion factors. The factor $\ell_{V}/\ell$ is obtained from the results for $a_{\rm lat}/\ell$ from our combined analysis in Section III using Eq. (39). The parameters of the dark energy model can then be constructed from the scale factor $A(\tau)$ and its derivatives with respect to $\tau$, including the Hubble rate $H=\dot{A}/A$, and the quantities $S=\ddot{A}/A$ and $T=\dddot{A}/\dot{A}$, where all three of these are measured in dual lattice units.
@@
and the derivatives are all taken under a jackknife to correctly propagate the errors through to the next stage of the analysis. As an example, the first derivative of the scale factor with respect to Euclidean time is
$\dot{A}(\tau/\ell)$ $\approx\frac{-A(\tau/\ell+2)+8A(\tau/\ell+1)-8A(\tau/\ell-1)+A(\tau/\ell-2)}{12},$
where the explicit factors of $\ell$ show that $\tau$ is measured in terms of renormalized dual-lattice ($\ell$) units.
@@
$3\left(-\frac{\dot{A}^{2}}{A^{2}}+\frac{\ell^{2}}{\ell_{V}^{2}}\frac{\ell_{V}^{2}}{A^{2}}\right)=\Lambda(H)\ell^{2}=(I_{H^{2}}+3\nu^{\prime}H^{2})\ell^{2}.$
The sign difference between the two terms on the left side of Eq. (IV.2) leads to a substantial cancellation between these terms. In order to construct this combination with sufficient precision to extract $\nu^{\prime}$, we require a precise determination of $\ell_{V}/\ell$ (which itself depends on having a precise $a_{\rm lat}/\ell$). This partial cancellation sets the precision level needed for the $a_{\rm lat}/\ell$ determination of Sect. III, which must be at the percent level in order to resolve a $\nu^{\prime}$ statistically and systematically different from zero. The value of $\nu^{\prime}$ is obtained from the slope of a linear fit to the left side of Eq. (IV.2) as a function of $H^{2}$, and $I_{H^{2}}\equiv\Lambda_{0}+3\tilde{\nu}S$ is obtained from the intercept. As we show in the following subsections, $\tilde{\nu}$ is consistent with zero. Taking $\tilde{\nu}=0$, we have $I_{H^{2}}=\Lambda_{0}$, and this becomes the input to our Method 2 determination of $a_{\rm lat}/\ell$, as detailed in Sect. III.3.2.
@@
In Sections IV.3, IV.4, and IV.5, most quantities ($\Lambda$, $H$, $\tau$, etc.) are evaluated in $\ell$ units. The exception to this is the four-volume, which is naturally expressed in units of $a_{\rm lat}^{4}$. Some fit functions in the following sections include a prefactor for the four-volume, which should be assumed to carry the correct units to convert between $a_{\rm lat}$ and $\ell$ where necessary. In these sections we suppress explicit factors of $\ell$ for clarity of the presentation.
@@ IV.3.1 Fits to $\Lambda(H)$ versus $H^{2}$
As discussed in the previous subsection, the value of $\nu^{\prime}$ is obtained by using the lattice shelling function to construct the left side of Eq. (IV.2). This leads to an effective cosmological constant $\Lambda(H)$, which must be linear in $H^{2}$ if it is to be consistent with Eq. (45). Fig. 13 shows the effective $\Lambda$ plotted versus $H^{2}$ for representative ensembles across four lattice spacings, where we find solid evidence that $\Lambda(H)$ is linear in $H^{2}$. The linear fit region that we choose is marked with black points in Fig. 13.
@@
Figure 13: $\Lambda$ versus $H^{2}$ plotted for four lattice spacings, along with a linear fit to the form Eq. (63). The points included in the fits are highlighted in black. Top left panel: $N_{4}=6$k, $\kappa_{2}=2.45$. Top right panel: $N_{4}=16$k, $\kappa_{2}=3.0$. Bottom left panel: $N_{4}=16$k, $\kappa_{2}=3.4$. Bottom right panel: $N_{4}=16$k, $\kappa_{2}=3.8$, $\beta=0.92$. The inflection point discussed in III.3.1 and the point corrresponding to the peak of the shelling function are highlighted in red.
@@ IV.3.2 Fits to $\nu^{\prime}$ as a function of four-volume
Our synthetic data sample is obtained from 10,000 random Gaussian draws for every ensemble, with a mean given by the central value of $\nu^{\prime}$ and the error given by the statistical and fit systematic error added in quadrature. We account for the error coming from $a_{\rm lat}/\ell$ to the arrays as follows: We assume that when $a_{\rm lat}/\ell$ for a given lattice spacing deviates from its central value by some percentage of one $\sigma_{a_{\rm lat}/\ell}$, all $\nu^{\prime}$ values at the same lattice spacing also shift by the same percentage of its $a_{\rm lat}/\ell$ systematic error, $\sigma_{\nu^{\prime},a_{\rm lat}/\ell}$. Therefore, for each lattice spacing, we generate an array of $10,000$ samples from a Gaussian random number generator, $A_{\rm ran}\thicksim\mathcal{N}(0,1)$. This allows us to construct the array $\left(\frac{a_{\rm lat}}{\ell}\right)_{\rm avg}+A_{\rm ran}\sigma_{\frac{a_{\rm lat}}{\ell}}$, which gives a distribution for $\frac{a_{\rm lat}}{\ell}$ with the correct central value and standard deviation at each lattice spacing. To obtain the final correlated array for $\nu^{\prime}$ we use the fact that for two independent normal distributions, $\mathcal{N}_{1}(\mu_{1},\sigma_{1})$, $\mathcal{N}_{2}(\mu_{2},\sigma_{2})$, their sum $\mathcal{N}_{1}+\mathcal{N}_{2}=\mathcal{N}(\mu_{1}+\mu_{2},\sqrt{\sigma_{1}^{2}+\sigma_{2}^{2}})$. We then get $\nu^{\prime}$ for each ensemble by subtracting $A_{\rm ran}\sigma_{\nu^{\prime},\frac{a_{\rm lat}}{\ell}}$ from the original $\nu^{\prime}$ sample, thereby including the $a_{\rm lat}/\ell$ systematic error in quadrature with the statistical and fit systematic error in $\nu^{\prime}$, and maintaining the correlation with $a_{\rm lat}/\ell$. The minus sign is due to the fact that the larger $a_{\rm lat}/\ell$ is, the smaller $\nu^{\prime}$ is. Because the random array for $a_{\rm lat}/\ell$ is shared by all ensembles at the same lattice spacing, the $a_{\rm lat}/\ell$ error is fully correlated across different volumes at a single lattice spacing. The fully correlated $a_{\rm lat}/\ell$ error covariance matrix is added to the diagonal matrix containing the statistical and fitting errors of $\nu^{\prime}$ in quadrature. This gives the full covariance matrix for $\nu^{\prime}$, which is reused when performing the fits over a jackknife array of $10,000$ random samples.
@@ IV.5 Alternative calculation of $\nu^{\prime}$
Unfortunately, the third derivative $\dddot{A}$ is only poorly determined by the data, given the large statistical errors associated with numerical derivatives of high order. There is also expected to be some cancellation between the $\tilde{\nu}$ term proportional to $S$ with that proportional to $T$, since $S$ and $T$ are equal up to corrections of ${\cal O}(H^{2})$. Thus, a meaningful constraint on $\tilde{\nu}$ is not possible with this approach.